OpenStudy (anonymous):

integrate sin³ 4x dx

OpenStudy (amistre64):

whenever you have an odd degree trig like that, its prolly a good idea to split it apart and swap about some identities

OpenStudy (amistre64):

or there is a formula that is a bit lengthy that you can try to recall.

OpenStudy (shubhamsrg):

simply let cos4x =t see if it helps ?

OpenStudy (anonymous):

no can't use the substitution method, because the derivative of the function is not available, what you need to do is expand the trig function using trig identities

OpenStudy (shubhamsrg):

Of course derivative of the function is available! What are you saying! :O

OpenStudy (anonymous):

you are saying t=cos(4x), then dt/dx=-4sin(4x), y=int t^4 so where are you leading this problem into more mess...

OpenStudy (anonymous):

you should assume t=4x i think this is what you meant...

OpenStudy (anonymous):

if this is the case then you can solve it using the integral formula for powers of sin or cos

OpenStudy (anonymous):

or we split that and use integration by parts after applying trig identities which will lead to the derivation of the reduction integral formula of trig to the power functions...

OpenStudy (shubhamsrg):

cos4x = t -4 sin4x dx = dt our ques is sin^3 (4x) dx= sin(4x)* sin^2(4x) dx = ((sin4x)dx)*(1-cos^2(4x)dx) = (1-t^2) dt/(-4) what part of this seems complicated to you?

OpenStudy (anonymous):

so you are integrating 1-t^2/(-4)

OpenStudy (anonymous):

integrate that you won't get the right answer...

OpenStudy (shubhamsrg):

-_-

OpenStudy (shubhamsrg):

Any reason why not?

OpenStudy (anonymous):

because when you integrate the function it contradicts with the reduction forumal answer

OpenStudy (anonymous):

even by using trig identity it can't be reduced to what you are kindly suggesting...

OpenStudy (anonymous):

i don't mind if there are different ways to solve the same problem but we should be able to justify our answers...

OpenStudy (anonymous):

split that trig function and using integration by parts, no way by any means that would be the same as 1-t^2

OpenStudy (anonymous):

let me give you a slang hint, why would we need a reduction formula, if it could be just solve by substituting the whole trig function...

OpenStudy (shubhamsrg):

You are making absurd hypothesis. How does it matter which path we chose when the final destination is the same! Of course the 2 answers will match whatever the path taken be, given that we have done it legitimately . Well mine is surely correct, check yours.

OpenStudy (anonymous):

this is my point i said they don't end up to the same results even by any trig expansion or simplification...

OpenStudy (anonymous):

the only way you can use u=cos(4x), after you use trig identity, not the other way round...

OpenStudy (anonymous):

you have to expand the rig identity in the form S^(n+1) then use the substitution method u=cos(4x)...

OpenStudy (anonymous):

so you need to make one substitution in order to apply u=cos(t),

OpenStudy (anonymous):

where t = 4x, otherwise it won't be valid... logically, unless you want to say you know it by head, but the student does not know

OpenStudy (anonymous):

now with those two substitution we can match with the reduction formula with no problem

OpenStudy (anonymous):

what i would use to save time on this would be Euler's identity.. for complex exponential and trig functions...

OpenStudy (anonymous):

in other words do not jump steps...that will reduce your mark